OpenStudy (pinklion23):

Hello open study people I was wondering if somebody could help me for well some time in math questions from 8th grade k12 :)

OpenStudy (pinklion23):

Choose the correct answer. Determine whether the given has a positive slope, a negative slope, a zero slope, or an undefined slope. https://static.k12.com/calms_media/media/213000_213500/213157/2/187f1c59866bfb925be32eadc01bb7b2ef3657e5/VHS_ALG_L2_S1_07_04_LQ1.gif A.undefined slope B.positive slope C.negative slope D.zero slope

OpenStudy (pinklion23):

@ganeshie8

OpenStudy (pinklion23):

A is what i think

OpenStudy (trojanpoem):

Horizontal or in other words parallel to the x axis , Slope = 0 Vertical or in other words parallel to the y axis , Slope = undefined

OpenStudy (pinklion23):

yeah thx

OpenStudy (pinklion23):

Which describes the slope of this line? https://static.k12.com/calms_media/media/1312500_1313000/1312636/1/b178d92bdc6205943ae2e5057d7f00724d0c718d/HS_A1G2_S1_07_04_ART4.jpg A.undefined slope B.zero slope C.negative slope D.positive slope

OpenStudy (trojanpoem):

Re-check the last comment, undefnied.

OpenStudy (pinklion23):

i know but there was a Zero Slope so.......

OpenStudy (pinklion23):

What is the slope of the line passing through the points (1, –5) and (4, 1)?

OpenStudy (pinklion23):

\[A.2 B. -\frac{ 4 }{ 5 }C.\frac{ 4 }{ 5 }D.-2\]

OpenStudy (pinklion23):

it will not let me put spaces in between them

OpenStudy (pinklion23):

@ParthKohli

OpenStudy (trojanpoem):

In the last question , to easily remember which is 0 / undefined use y_2 - y_1/ x_2 - x_1 in case of the second line we have x as a number between 1, 2 assume it's 1.5 so y_2 - y_1/ 1.5 - 1.5 = y_2 - y_1/ 0 ( denominator is zero so undefined) in case of horizontal line the y will be the same so 0/x_2 - x_1

OpenStudy (trojanpoem):

What is A, B, C ,D ? matrices ?

OpenStudy (trojanpoem):

are

OpenStudy (pinklion23):

wait so in number 2 the answer is undefined slope right?

OpenStudy (trojanpoem):

For the slope of the line passing throught two points : \[(x_{1}, y_{1}) (x_{2}, y_{2}) \] This formula is generally used : \[\frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\]

OpenStudy (trojanpoem):

Yeah, I was explaining how I recall them easily.

OpenStudy (trojanpoem):

"Re-check the last comment, undefnied." I didn't mean I was mistaken , I meant I answered the question of yours pre-posting it.

OpenStudy (pinklion23):

great so number 3 the answer is..... hold on

OpenStudy (pinklion23):

i got 9?

OpenStudy (pinklion23):

er 0/9

OpenStudy (trojanpoem):

(1, -5) ( 4, 1) (1 - (-5)) / (4 - 1) = (1+5)/ 3 = 6/3 = 2

OpenStudy (pinklion23):

I'm sorry i just am not in the zone :'(

OpenStudy (pinklion23):

What is the slope of the line passing through the points (0, 4) and (−8, −1)?

OpenStudy (pinklion23):

@phi

OpenStudy (trojanpoem):

-1 - 4/ -8 - 0 = -5/-8 = 5/8 , mate it's direct substitution

OpenStudy (pinklion23):

sorry

OpenStudy (pinklion23):

What is the slope of the line whose equation is 12 = 4x – 6y?

OpenStudy (pinklion23):

if i am annoying you, you do not have to help :/

OpenStudy (pinklion23):

phi can you help me?

OpenStudy (phi):

***slope of the line whose equation is 12 = 4x – 6y?*** to find the slope, you want to make the equation look like y = m x + b (that is called slope-intercept form). Have you see this before?

OpenStudy (pinklion23):

yes

OpenStudy (phi):

if you start with 12 = 4x -6y and add 6y to both sides you get 6y+12 = 4x -6y + 6y do you know how to simplify the right-hand side? i.e. what is -6y+6y ?

OpenStudy (pinklion23):

0

OpenStudy (pinklion23):

so then you have 6y+12=4x

OpenStudy (phi):

yes. now add -12 to both sides

OpenStudy (pinklion23):

um i can't because then it would look like tis...... 67=4x+12

OpenStudy (pinklion23):

6y sorry

OpenStudy (phi):

6y+12=4x 6y+12 - 12 = 4x -12 the 12-12 is 0 on the left you get 6y= 4x-12

OpenStudy (pinklion23):

right

OpenStudy (pinklion23):

but i can't do 4x-12 because 12 does not have and x

OpenStudy (phi):

you can add (or subtract) numbers from letters 4x-12 is OK what you can't do is "simplify" that, the same way you could simplify 4x-12x now divide both sides by 6 \[ \frac{6y}{6}= \frac{4x-12}{6} \\ y= \frac{4}{6}x - \frac{12}{6} \]

OpenStudy (pinklion23):

oh yeah i remember now wow thx so the answer is 2? but i can not put in 2

OpenStudy (pinklion23):

i have to put in....... oh wait it is a rid sorry yeah so 12/6 is the answer

OpenStudy (pinklion23):

grid*

OpenStudy (phi):

if you simplify \[ y= \frac{4}{6}x - \frac{12}{6} \] you get \[ y= \frac{2}{3} x -2\] the slope is the number in front of the "x"

OpenStudy (pinklion23):

so the answer to the question is 2/3?

OpenStudy (phi):

yes. once you have y= m x + b the m (number in front of the x) is the slope and b is the "y-intercept"

OpenStudy (pinklion23):

great thx so 2/3 is the answer for number 5 thx a bunch can you help me more?

OpenStudy (phi):

if you make a new post.

OpenStudy (pinklion23):

yes i will :)